JOURNAL OF GEOPHYSICAL RESEARCH, VOL. ???, XXXX, DOI:10.1029/,

1 Validation of Ozone Monitoring Instrument Nitrogen 2 Dioxide Columns

E. A. Celarier,^{1} E. J. Brinksma,^{2} J. F. Gleason,^{3} J. P. Veefkind,^{2} A. Cede,^{4} J. R. Herman,^{3} D. Ionov,^{5,6} F. Goutail,^{6} J-P. Pommereau,^{6} J-C. Lambert,^{7} M. van Roozendael,^{7} G. Pinardi,^{7} F. Wittrock,^{11} A. Schönhardt,^{11} A. Richter,^{11} O. W. Ibrahim,^{13} T. Wagner,^{12} B. Bojkov,^{4} G. Mount,^{8} E. Spinei,^{8} C. M. Chen,^{9} T. J. Pongetti,^{9} S. P. Sander,^{9} E. J. Bucsela,^{4} M. O. Wenig,^{4} D. P. J. Swart,^{10} H. Volten,^{2,10} M. Kroon,^{2} and P. F. Levelt^{2}

Edward A. Celarier, SGT, Inc. 7701 Greenbelt Rd, Greenbelt, MD 20770 (Edward.Celarier@gsfc.nasa.gov)

Ellen Brinksma, KNMI, 3730 AE De Bilt, The Netherlands (ellen.brinksma@knmi.nl)

1SGT, Inc., 7701 Greenbelt Rd Ste 400,

X - 2 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

3 Abstract. We review the standard nitrogen dioxide (NO_{2}) data prod-4 uct (Version 1.0.0), which is based on measurements made in the spectral 5 region 415–465 nm by the Ozone Monitoring Instrument (omi) on the EOS-6 Aura satellite. A number of ground- and aircraft-based measurements have 7 been used to validate the data product’s three principal quantities: strato-8 spheric, tropospheric, and total NO_{2} column densities. It is found that omi 9 underestimates the stratospheric column by about 14%, and underestimates 10 both the total and tropospheric columns by 15 − 30%. The validation of 11 the omi NO_{2} data is complicated by the fact that some of the ground based 12 measurements have not been well validated, themselves, and by problems in-13 herent in comparing point measurements from the ground to large area av-14 eraged satellite measurements. In addition, some of the ground based mea-15 surements have only been made over a relatively short period of time, lim-16 iting their use for statistical comparison with omi. Nonetheless, by using all 17 the available ground based measurements, it is possible to arrive at some quan-

titative characterization of the biases in the omi NO_{2} data product.

18

Greenbelt MD 20770, USA.

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 3

1. Introduction 19 The Ozone Monitoring Instrument (omi) is a space borne spectroradiometer that uses 20 a 2-dimensional ccd array detector to simultaneously measure the spectra of the earth 21 shine radiance at large number of viewing angles, approximately transverse to the Aura 22 spacecraft’s flight track. omi measures in three broad spectral regions (uv-1, uv-2, and 23 vis), with a spectral resolution on the order of 0.5 nm. Applying spectral fitting techniques 24 to the omi data permits the simultaneous retrieval of a wide range of atmospheric trace 25 gas concentrations as well as cloud and aerosol properties and loadings. Among the trace 26 gases that can be retrieved, ozone (O_{3}) and nitrogen dioxide are identified as essential 27 measurements, both for the ongoing monitoring of the Earth’s stratospheric ozone layer 28 and for the monitoring of tropospheric air quality. A more extensive discussion of the omi

instrument itself can be found in Levelt and Bhartia [2007].

29

30 The omi NO_{2} data production algorithm is designed to retrieve total vertical column 31 densities of NO_{2} and separate stratospheric and tropospheric column densities; this en-32 ables the improvement in the calculation of the total vertical column. This separation is 33 important (and possible) because the chemistry and dynamics of NO_{2} are different be-34 tween the stratosphere and the troposphere. Accurate measurements of the tropospheric 35 NO_{2} are significant for the characterization of air quality, a primary objective of the Aura 36 and omi missions. 37 This paper discusses only the validation of the omi NO_{2} standard product, archived at 38 the Goddard Earth Sciences Data and Information Services Center (ges-disc). We do not

2Royal Netherlands Meteorological

X - 4 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

39 discuss any other product, such as the Level-4 product, archived at the Royal Netherlands 40 Meteorological Institute (KNMI). Furthermore, comparison of omi to other spaceborne 41 NO_{2} sensors, which all have different equator crossing times from omi is beyond the scope

of this paper.

42

43 A number of efforts toward NO_{2} validation have been initiated, in which measurements 44 are made coincident with omi overpass measurements. The purpose of this paper is to 45 provide an overview of results from these efforts. A variety of instruments and techniques 46 have been used, each with its characteristic sensitivity to stratospheric, tropospheric, 47 or total column NO_{2}. This paper will address the advantages and, in some cases, the 48 limitations of the various measurements. 49 Several of the techniques described are new, and have not been well-validated. In addi-50 tion, there is an essential difference between observations of NO_{2} taken from the ground, 51 and observations averaged over a satellite field of view (fov). Spatial inhomogeneity, 52 characteristic of airborne constituents emitted at (possibly moving) point sources, and 53 subject to surface-level winds, implies that a single point measurement will often not be 54 a representative sample within a “collocated” satellite fov covering a region of the order 55 of several hundred square kilometers. Monthly average comparisons of ground-based and 56 satellite measurements can remove much of the variability due to fov-point measurement 57 differences. A preliminary measurement of horizontal inhomogeneity in the NO_{2} field is 58 presented in the dandelions overview paper [Brinksma et al., 2007], using a set of si-59 multaneous tropospheric NO_{2} measurements made at different azimuths. Veefkind et al. 60 [2007] show a comparison of regridded omi NO_{2} data with ground-based observations by

Institute, 3730 AE De Bilt, The

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 5

61 the Dutch national air quality network. This network distinguishes regional stations, and 62 city and street stations, which are close to source regions. For the period of the satellite 63 data, NO_{2} reported by 35 stations around the Netherlands, averaged between 11:00 and 64 14:00 h local time, was compared to the collocated omi measurements. A strong correla-65 tion (R = 0.94) between the satellite data and the regional station data was found. By 66 contrast, correlations with urban stations are weak, because local conditions may vary 67 strongly over a few hundred meters, far smaller than the spatial resolution of omi. Be-68 cause of both the novelty of some of the techniques, and such spatial-scale effects, specific

results are sometimes limited to qualitative, or order-of-magnitude, conclusions.

69

70 Among the validation studies discussed here are ground based observations made within 71 the saoz and doas networks. These are zenith-sky, twilight measurements, which are 72 sensitive to the stratospheric NO_{2} column (as explained in section 4.1.1). In addition, 73 the locations are very often pristine areas, or at elevation. Even if the instrumentation 74 were sensitive to tropospheric NO_{2}, the lack of pollution would lead to measurements

dominated by the stratospheric NO_{2} amount.

75

76 To focus on polluted areas, where satellite NO_{2} retrievals are most challenging, novel,

or as yet unvalidated techniques must be used.

77

78 Measurements of scattered light by the MultiAxial doas (max-doas) technique [Platt, 79 1994; Wittrock et al., 2004], using a range of viewing angles, from nearly horizontal through 80 zenith, are sensitive to the tropospheric part of the column, and provide both total and

Netherlands.

X - 6 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

81 tropospheric NO_{2} amounts. We present results from max-doas measurements taken in

a polluted area, but away from immediate local sources, in section 4.2.1.

82

83 Direct-sun ground based measurements, made with a Brewer spectrophotometer [Cede 84 et al., 2006] and with newly-developed direct-sun instruments, including a high-resolution 85 Fourier-transform uv-fts technique [Cageao et al., 2001], and a number of direct-sun 86 doas-type measurements, are sensitive to the total NO_{2} column. We will briefly review

some preliminary results from these methods in section 4.3.

87

88 Validation of the omi NO_{2} data should take account of the sensitivity of the numerous 89 geophysical and geometric algorithmic inputs. These include the a priori profile shapes, 90 surface albedo, and measured and assumed cloud properties. These, in particular, greatly 91 affect the air mass factors (amf; the ratio of slant-column density of the absorber along 92 the optical path to the vertical column density) the algorithm calculates. Tropospheric 93 NO_{2} profiles have been measured with lidar in the Netherlands, during a number of days in

September 2006, and with airborne instrumentation during various validation campaigns.

94

95 During the Polar Aura Validation Experiment (pave) (flights from New Hampshire, 96 January and February 2005) the td-lif instrument [Thornton et al., 2000; Cleary et al., 97 2002] was used for in situ sampling of NO_{2}, during the aircraft flights. The NASA DC-8 98 performed two flight legs at 300 m altitude, near the top of the boundary layer. When

flights entered the boundary layer, strongly enhanced concentrations of NO_{2} were found.

99

100 During the intex-b campaign (flights from Houston, Texas, March 2006 and from 101 Honolulu, Hawaii, and Anchorage, Alaska, April and May 2006) the td-lif instrument

3Code 613.3, NASA Goddard Space

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 7

102 measured NO_{2} in situ. Spirals were flown by the NASA DC-8 during several flights in

spatial and temporal collocation with omi observations.

103

104 Besides intex-b, a small number of other airborne campaigns have been carried out, 105 measuring NO_{2} in situ, and have been applied to satellite validation [Heland et al., 2002;

Martin et al., 2006].

106

107 The remainder of Section 1 contains a brief description of NO_{2} chemistry in the strato-108 sphere and troposphere, the omi measurement of NO_{2}, and the availability of the data 109 sets. Section 2 describes the omi measurement, and Section 3 describes the algorithm that 110 reduces the raw omi measurements to the various NO_{2} columns. The various subsections 111 of Section 4 discuss the validations of the three principal NO_{2} products: stratospheric 112 column (Section 4.1), tropospheric column (Section 4.2), and total column (Section 4.3).

The conclusions are presented in Section 5.

113

1.1. Nitrogen dioxide in the stratosphere 114 Nitrogen dioxide participates both directly and indirectly in the catalytic destruction of 115 ozone in the stratosphere [Finlayson-Pitts and Pitts, 2000]. Direct catalytic destruction 116 takes place via a NO − NO_{2} cycle that catalyzes the reaction O_{3} + O → 2O_{2}, while 117 while NO_{2} concentrations indirectly control ozone loss through other catalytic cycles by 118 controlling, for example, the distribution of chlorine between its catalytically active (ClO) 119 and inactive, reservoir (ClONO_{2}) species. In the stratosphere, NO_{2} concentration has a 120 distinctive diurnal cycle, due largely to the photochemistry of nitrogen oxides. At night, 121 all the photolysis reactions stop, shifting the steady state to NO_{2}. NO_{2} is converted,

Flight Center, Greenbelt MD 20771, USA.

X - 8 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

122 through ozonolysis, to NO_{3}, which can further combine with NO_{2} to form N_{2}O_{5}. This 123 results in a slow decrease of NO_{2} over the course of the night. When the air is again 124 sunlit, the N_{2}O_{5} rapidly redissociates to NO_{2} and NO_{3}, which photolyzes instantaneously, 125 mostly to NO. Meanwhile, NO_{2} photolyzes very rapidly, and so decreases very rapidly 126 at sunrise. While in daylight, the dominant processes are the interconversion between 127 NO and NO_{2}. Under most typical conditions of temperature and ozone concentration, 128 latitude, and season, there is a slow increase in NO_{2} concentration over the course of the 129 daylight hours. In addition to the chemical and photochemical processes, transport by 130 the winds, particularly in the vicinity of the polar jets may mean that the air that one is 131 measuring has not had the photochemical history one would expect, based on location and 132 local time, alone. Some caution is therefore needed in matching satellite measurements

to the ground based measurements.

133

1.2. Nitrogen dioxide in the troposphere 134 In the troposphere, nitrogen oxides are a significant contributor to poor air quality. 135 Gaseous NO_{2} is red in color, and gives rise to the characteristic brownish cast of polluted 136 air. Both NO and NO_{2} are harmful to lung tissue, and, as a powerful oxidizing agent, 137 NO_{2} is harmful to biological tissue generally. Besides its direct effects, photolysis of NO_{2}

contributes to ozone production [Finlayson-Pitts and Pitts, 2000] according to

138

NO_{2} + hν ⟶ NO + O
O + O_{2} ⟶ O_{3}

4GEST Program, University of Maryland

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 9

139 Nitrogen oxides are produced in high-temperature processes, most notably in com-140 bustion (fossil fuels and biomass burning) and in lightning. As a rule, the higher the 141 combustion temperature, the more NOx is produced. Nearly all the NOx (NO + NO_{2}) 142 that is significant for human health is produced by industrial and urban activity, including

transportation and power generation.

143

144 NO_{2} is removed from the troposphere through conversion to HNO_{3}, nitric acid, which 145 readily dissolves in any available water droplets. NO_{2} plumes are detected only up to 146 about 200 km from their source. In the neighborhood of industrial or urban sources, 147 there is a distinct diurnal pattern in the concentration of NOx. The diurnal signal at any 148 location is the result of a complex interplay between the emission source field in space 149 and time, photochemical lifetimes, advection, and the concentrations of chemical sinks 150 for NOx species. As mentioned before, these also give rise to spatial inhomogeneities on 151 a sub-100 m scale. At mid- to low-latitudes, where a polar-orbiting satellite passes over 152 a given location is only once or twice a day, the satellite only sees a “snapshot” of the 153 state of the polluted atmosphere at the overpass times. In the mid-to-upper latitudes, 154 inconsistent measurements, from one orbit to the next, over some location, may well result 155 from significant changes in the NO_{2} concentrations over the intervening 100 minutes, as

well as from other rapid geophysical changes, e.g. in cloud cover.

156

1.3. OMI measurement of NO_{2} 157 The Aura satellite is a polar-orbiting, sun-synchronous satellite, whose orbital period 158 is 99 minutes. Aura flies over the entire surface of the Earth every 14–15 orbits. Using

Baltimore County, Baltimore MD, USA.

X - 10 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

159 the 2-dimensional ccd array detector, with pixel binning factors chosen to optimize the 160 signal-to-noise ratio, the instrument measures earthshine radiance spectra simultaneously 161 in 60 effective fields of view (fov), approximately transverse to the flight track, every 2 162 seconds (the ccd is read out every 0.4 s, and co-added in groups of 5), over a range of 163 angles 57^{∘} either side of nadir. This gives a sufficient “push-broom” width to view the 164 entire sunlit surface of the Earth, even in the tropics, with multiple orbital overlaps for

much of the mid- to high-latitude regions.

165

166 During normal operations, omi measures the solar irradiance spectrum once every 24 167 hours. The ratio of the earthshine radiance to solar irradiance, the spectral albedo, is

calculated for each fov. The omi NO_{2} retrieval algorithm is described in Section 3.

168

1.4. Data availability

The omi NO_{2} data product is available in a number of different geospatial forms:

169

• Level-2 orbital swath (L2);

170

• Daily global gridded, 0.25^{∘} × 0.25^{∘} (L2G);

171

• Station and regional overpass (OVP).

172

173 The L2 and L2G datasets and associated documentation are freely available through the 174 NASA Goddard Earth Sciences Data and Information Services Center (ges-disc, url:

   http://daac.gsfc.nasa.gov/Aura/OMI/index.shtml).

175

176 The ovp data, generated daily for over 100 locations around the world, and also in 177 support of validation and regional pollution studies, are available through the Aura Vali-

178 dation Data Center (avdc, url: http://avdc.gsfc.nasa.gov/Data/Aura/OMI/OMNO2/

5Department of Atmospheric Physics,

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 11

179 index.html). The subsetted data used for this paper were generated at the avdc using

the recommended usage quality flags [Celarier et al., 2006].

180

181 Both the L2G and ovp data products are derived from the L2 data set, and not all 182 of the fields found in the L2 data may be found in the derived data products. Complete 183 details concerning the contents of the Level-2 files are available in Veefkind and Celarier

[2006].

184

185 Each L2G file contains a 0.25^{∘} ×0.25^{∘} grid data structure. Each cell of the grid contains 186 a stack of data values for all the fovs whose centers fell within that cell. For each fov a 187 subset of the available L2 fields is stored. Because it is organized geographically, the L2G 188 data set should be suitable for users who wish to study specific geographic locations, as, 189 for example, in the case of validation against ground based measurements, or for regional

air quality studies.

190

191 Though all the data have been publicly released and are freely available, prospective 192 data users are strongly encouraged to contact the principal investigators responsible for 193 the data sets.

2. OMI Measurement 194 Atmospheric NO_{2} column densities are retrieved using spectral measurements of the 195 solar irradiance and earth shine radiance in the wavelength region 415–465 nm, using 196 the instrument’s VIS detector. The measurements are made with a spectral resolution of 197 ∼ 0.5 nm. Daily measurements of the solar irradiance have been made since the instru-198 ment became operational, with the exception of the period 2006 February 28 through 2006

Research Institute of Physics, St.

X - 12 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

199 March 3, when a problem with the instrument’s folding mirror prevented making daily 200 irradiance measurements. Using measured irradiance spectra has resulted in the appear-201 ance of stripe structure in virtually all the data products, in which the retrieved quantities 202 have different means at each of the 60 cross-track positions. This has necessitated the

implementation of “destriping” algorithms.

203

204 The omi instrument design and performance have been described by Levelt et al. 205 [2006b, a]. Dobber et al. [2006] have discussed the calibration of the instrument, and

the origin of the striping, or cross-track bias.

206

3. OMI Algorithm 207 In this section we present the essential details of the algorithm. A much more detailed 208 description of the omi NO_{2} algorithm, its theoretical underpinnings, and uncertainty 209 analysis may be found in [Bucsela et al., 2006; Boersma et al., 2002; Wenig et al., 2007;

Boersma et al., 2004].

210

3.1. Slant column densities 211 The first part of the calculation of NO_{2} columns consists in calculating the slant column 212 densities. Since the omi-measured radiance and irradiance spectra, and the laboratory 213 spectra are all measured on different wavelength scales, the measured spectra are in-214 terpolated onto a common scale. The spectral albedo, R, is then fitted by a nonlinear

least-squares technique onto the function

215

R(λ) = P_{3}(λ) · exp (−c_{NO2} · σ_{NO2}(λ) − c_{O3} · σ_{O3}(λ)) ·

Petersburg State University, St. Petersburg,

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 13

(1 + c_{ring} · σ_{ring}(λ)) ,  (1)

216 where σ is the absorption cross section of the indicated species, and P_{3} is a third-order 217 polynomial in the wavelength, which models the component of the spectrum that is 218 smoothly varying, due to Rayleigh and Mie scattering. Literature spectra are used for 219 σ_{NO2} [Vandaele et al., 1998], σ_{O3} [Burrows et al., 1999a], and σ_{ring} [Chance and Spurr, 220 1997]. These spectra were convolved with a model omi instrument slit function prior to 221 use in the fitting algorithm. In all, each measured spectrum is subjected to a nonlin-222 ear least-squares fit with a total of seven free parameters (c_{NO2}, c_{O3}, c_{ring}, and the four 223 coefficients in P_{3}(λ)). The algorithm also estimates the uncertainties in each of the fit

parameters, as well as the χ^{2} error and R.M.S. error of the fit.

224

225 The determination of the scd does not include fitting to the absorption spectra of either 226 H_{2}O or O_{2} − O_{2}. The introduction of these species were found not to have a large effect

on the retrieved NO_{2} scd, and resulted in longer execution times.

227

3.2. Initial vertical column densities 228 Initial estimates of the vertical column density (V_{init}) are calculated using amfs de-229 rived from typical climatological profile shapes, with a nominal amount of NO_{2} assumed 230 in the troposphere (amf_{init}). That is, the initial vertical columns are computed under 231 the assumption that the troposphere is not polluted. The profiles assumed are a 1-year

Russia

X - 14 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

232 average of daily profiles computed using the geos-chem model in the troposphere, and

the Goddard Chemical Transport Model in the stratosphere [Bucsela et al., 2006].

233

3.3. Stratosphere-troposphere separation 234 At the core of the omi NO_{2} algorithm is a procedure to identify fields-of-view (fov) 235 where there is significant tropospheric NO_{2}. This is required because the air mass factor 236 depends upon the profile shape (though not the total amount, since the trace gas is op-237 tically thin): fovs where there is significant tropospheric NO_{2} require a different amf to 238 compute the vcd from the scd. It is observed [Gordley et al., 1996] that the stratospheric 239 NO_{2} field has relatively small gradients, particularly in the zonal direction. Our procedure 240 for the stratosphere-troposphere separation essentially identifies the slowly-varying com-241 ponent of the total NO_{2} field as the stratospheric field, and the rest as the tropospheric 242 field. 243 Each orbit is treated as follows. The “target” orbit’s data are read in, along with the 244 data from all other available orbits that were measured within ±12 hours of the target. 245 Each fov is identified with a grid cell on a 1^{∘} × 1^{∘} grid in latitude and longitude. For 246 all the fovs that are identified with a particular grid cell, a “cost” is computed from 247 the initial amf and uncertainty estimate for the V_{init}; the value of V_{init} having the lowest 248 cost is saved in its associated grid cell. A “mask” identifying grid cells where there are 249 known, persistent sources of NO_{2} was developed for use in the algorithm; no V_{init} values 250 are stored in masked grid cells. The V_{init} values are averaged in the meridional direction 251 with a boxcar function of half width 5^{∘}. For each 1^{∘} latitude band, a wave analysis is

6CNRS, Route des Gatines, 91370

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 15

252 performed, fitting waves 0, 1, and 2, to give a preliminary background field. Grid cells 253 whose V_{init} value exceeds the preliminary background field by more than one standard 254 deviation are then excluded, and the wave analysis is redone. The result of this is a 255 background field (V_{bg}) that has been influenced very little by the presence of regions 256 of high NO_{2} concentration. Since the V_{init} values were obtained using an amf that is 257 appropriate to a profile having most of the NO_{2} in the stratosphere, no further correction

to the background field is required.

258

3.4. Vertical column densities 259 For each fov, the value of V_{init} is compared to the evaluated background field at that 260 location. If V_{init} is less than the background field, then the final value of V (the total NO_{2} 261 column amount) is taken to be V_{init}. If V_{init} is larger than the background field, then the 262 “polluted” part (V_{init}−V_{bg}) is scaled by the ratio amf_{init}/amf_{pol}, where amf_{pol} is obtained 263 using the climatological geos-chem-modeled profile [Bey et al., 2001; Martin et al., 2002]. 264 This procedure gives the total column, the background column, and the polluted column. 265 In addition, a tropospheric column, equal, in the polluted case, to the polluted column 266 plus the amount of the unpolluted profile that exists below the tropopause (assumed to 267 be at 200 hPa). Finally, if, according to the standard cloud product, the cloud fraction is

Verrières le Buisson, France

X - 16 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

268 larger than 0.1, then the “below cloud amount” (the amount of NO_{2} that is inferred to

be below the visible surface of the clouds) is also computed.

269

3.5. Destriping 270 Due to radiometric calibration and dark-current drift in omi’s ccd detectors, which 271 affects the radiance measurements differently from the irradiance measurements, nearly 272 all omi Level-2 data products show some degree of cross-track bias, which appears as 273 stripes of systematically elevated or diminished values at certain cross-track scan positions, 274 persisting throughout each orbital track [Dobber et al., 2006]. While the origin of much 275 of the cross-track bias is now understood, and an improvement in the Level-0 to Level-1 276 processing algorithm is being implemented, the data available for the purpose of validation

to date have had significant cross-track bias.

277

278 A “destriping” procedure has been implemented in the omi NO_{2} algorithm. In this 279 procedure, the NO_{2} scds and amfs are collected for the 15 orbits (or fewer, depending on 280 data availability) used to construct the background field. These are then used to construct 281 separate scd correction offsets for the northern and southern hemispheres:

Δ_{i} = \overline{SCD_{i}} − \overline{AMF_{i}} · \begin{matrix}⟨ & SCD & ⟩ \\ ⟨ & AMF & ⟩ \end{matrix} ,  (2)

282 where i is the cross-track scan position (1 to 60), the overlines indicate averages for single 283 scan positions, and angle-brackets indicate averages over all scan positions. The Δ_{i} are

subtracted from the scds before applying the final air mass factors.

284

285 One concern about this procedure has been that it could introduce an unknowable bias 286 in the computed NO_{2} column densities. This will be discussed in light of the ground-based

7Chemistry and Physics of Atmospheres,

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 17

287 validation data. Another concern is that it is observed that, even using this procedure, 288 there is some residual striping in the data. The data that have severe bias are generally 289 flagged (and so not used in the validation). We assume that the residual biases are small, 290 and are sufficiently symmetrically distributed, that the biases cancel: The cross-track 291 biases change from orbit to orbit, and successive overpasses of any site are viewed with 292 different fovs.

4. Validation of OMI NO_{2} columns

4.1. Stratospheric column 293 4.1.1. SAOZ and DOAS instruments in the NDACC network 294 The NDACC (Network for the Detection of Atmospheric Composition Change) is an 295 international cooperative network that coordinates the operations and data analysis at 296 more than 30 stations at various latitudes on the globe from 76^{∘}S to 79^{∘}N. The ground-297 based UV-Visible zenith-sky spectrometers include both saoz (Système d’Analyse par 298 Observations Zénithales) as well as doas instruments, which provide ozone and NO_{2} ver-299 tical columns at sunrise and sunset using the Differential Optical Absorption Spectroscopy 300 (doas) technique [Platt, 1994] in the spectral range 410–530 nm. Zenith-sky measure-301 ments made at solar zenith angles between 86–91^{∘} are averaged to give estimates of the 302 column NO_{2}. Because of the optical geometry of the measurement, the retrieved NO_{2} 303 column is much more sensitive to the stratospheric NO_{2} column than to the tropospheric 304 column. Most of the instruments are located in remote geographical regions, far from 305 any significant source of tropospheric NO_{2}. Figure 1 shows the geographical distribution

Federal Space Pole, Belgian Institute for

X - 18 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

306 of the saoz stations that are operated by the French Centre National de la Recherche 307 Scientifique (CNRS). Only the instruments at the Observatoire Haute-Provence (OHP), 308 France, and Bauru, Brazil, are in any proximity to presumed anthropogenic sources of 309 NO_{2}. Measurements from the saoz instruments have been previously used to compare 310 with NO_{2} measurements by the space borne gome [Burrows et al., 1999b] and scia-311 machy [Bovensmann et al., 1999] instruments [Ionov et al., 2006a, b, 2007; Piters et al.,

2006; Lambert et al., 2001].

312

313 Stratospheric NO_{2} exhibits a pronounced diurnal cycle due to its daytime photolysis 314 into NO and nighttime conversion into N_{2}O_{5}. The NO_{2} daily cycle starts with a fast 315 drop shortly after sunrise, followed by a quasi-linear slow increase during the day, a fast 316 increase at sunset, and finally a slow decrease during the night. The time-dependence of 317 the stratospheric NO_{2} concentration has important implications for the validation of the 318 space based NO_{2} measurements. If the ground based measurements are not collocated 319 in time with the omi measurements, they need to be corrected, using photochemical and 320 transport models, to account for the time difference. In addition, if the ground-based 321 measurements entail an optical path that is more horizontal than vertical, view and solar

geometries must be taken into consideration when identifying “collocated” measurements.

322

323 The diurnal cycle has been simulated with a photochemical box model derived from 324 the slimcat 3d chemical-transport model [Denis, 2005]. It includes 98 chemical and 39 325 photochemical reactions, including heterogeneous chemistry on liquid and solid particles. 326 Calculations are made at 17 altitude levels with a time step of 1 minute. The NO_{2} total 327 column is obtained by integrating the profile assuming a constant density in each layer.

Space Aeronomy, 3 Avenue Circulaire,

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 19

328 Figure 2 shows the results of simulations at two saoz stations, OHP at mid-latitude and 329 and Scoresby Sund in the Arctic, for spring and fall. Using this photochemical model, a

B-1180 Brussels, Belgium

8Laboratory for Atmospheric Research, Dept. of Civil and Environmental Engineering, Washington State University, Pullman, WA 99164-2910, USA

9Science Division, NASA Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109

10Netherlands National Institute for Public Health and the Environment (RIVM), Bilthoven, The Netherlands

11Institute for Environmental Physics, University of Bremen, D-28334 Bremen, Germany

12Max-Planck-Institute for Chemistry, Mainz, Germany

13Institute for Environmental Physics, University of Heidelberg, D-69120 Heidelberg, Germany

X - 20 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

330 diurnal time series of the ratio NO_{2}(sunrise)/NO_{2}(t) was calculated for each month, and 331 at each saoz location. As saoz is an average of measurements between 86^{∘} and 91^{∘} SZA 332 the NO_{2} column at 88.5^{∘} sza is taken as the sunrise reference. The omi stratospheric 333 measurements (total minus tropospheric columns) were calculated and then normalized

to corresponding sunrise values using these ratios.

334

335 The optical geometry of the twilight saoz measurements is such that the light paths 336 traverse rather large distances through the stratosphere, so the stratosphere is sampled at 337 some distance from the measurement site. This should be taken into account when seeking 338 “match up” satellite fovs corresponding to the ground-based measurements, especially in 339 regions with large stratospheric NO_{2} gradients. However, in the present studies, a simpler 340 approach was adopted, in which the match up criterion was that the ground site was 341 within an OMI field-of-view. If multiple match ups are identified for a single day, the one

whose center fell closest to the ground station was used.

342

343 Finally, the omi NO_{2} algorithm provides total column NO_{2}, and the tropospheric col-344 umn NO_{2}. Since saoz measurements are roughly 50 times as sensitive to the stratospheric 345 column as to the tropospheric column, it is of interest to compare the saoz-derived values

to the difference of the total and the tropospheric columns.

346

347 Figure 3 shows the time series of the difference between the sunrise saoz measurements 348 and the matching omi measurements from eight saoz sites, adjusted to account for the 349 difference between the satellite overpass time and sunrise. The statistical characteristics of 350 these differences are presented in Table 1. Besides a comparison to just the stratospheric 351 column, the table presents a comparison between the omi total column and the saoz 352 instrument measurements. It is seen that, at virtually all latitudes, there is very good

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 21

353 agreement, on average, between the ground-based and satellite-based measurements of 354 the total stratospheric NO_{2} column. However, a small annual cycle is apparent in the 355 time-series for the higher latitudes, with lower values in the winter than in the summer. 356 This cycle, which appears in both the northern and southern hemisphere high latitudes, 357 may be related to the omi sampling under those conditions, or may reflect a sensitivity 358 to the choice of matching omi fov corresponding to a given ground-based observation, 359 or may be due to a bias either in the satellite measurement at high solar zenith angle, 360 or in the ground-based measurements as the sunrise azimuth tends poleward. While the 361 influence of the seasonal cycle on the overall statistics is fairly small, understanding it

may be an avenue of further study.

362

363 The correlation coefficients between the saoz and omi-measured stratospheric NO_{2} 364 columns are better, and the mean absolute differences smaller, for the mid-to-high latitude 365 sites than for the tropical sites. Since the stratospheric NO_{2} concentrations are smaller 366 in the tropics in the first place (annual mean of about 2.5 × 10^{15} cm^{−2}, compared to an 367 annual mean of 4 − 5 × 10^{15} cm^{−2} at the high latitude sites), the relative differences are 368 much greater in the tropics, and even the mid-latitude sites (OHP and Kerguelen), than

at the high-latitude sites.

369

4.2. Tropospheric column 370 4.2.1. MAX-DOAS 371 The MultiAxis doas (max-doas) technique is an extension of the zenith-sky doas 372 technique described in Section 4.1.1, but having much greater sensitivity to lower tro-373 pospheric layers. In brief, a max-doas typically consists of two main parts: A grating 374 spectrometer mounted inside a thermostatted box that is located inside a building, and

X - 22 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

375 one or more scanning telescopes connected to the spectrometer via fiber optics. Consec-376 utive measurements at increasing elevation angles are performed in an acquisition cycle

that always contains observations at a number of low elevations, and a zenith observation.

377

378 From each of the measurements, a slant column is retrieved using the DOAS method 379 described in Section 4.1.1[Platt, 1994]. Besides NO_{2}, a number of other absorbers, plus 380 the Ring effect, are included in the fit, as are a multiplicative polynomial and an additive 381 polynomial for stray light correction. In order to account for the temperature dependence 382 of the NO_{2} absorption spectrum, a second cross-section (295 K and 221 K) may be intro-383 duced in the retrieval to improve the fit and correct the derived vertical column. Some 384 groups use one-temperature retrievals, while other groups use two-temperatures, and this 385 may cause some discrepancy between the max-doas instrument results. In contrast, 386 the omi retrieval uses a single temperature, and applies a height-dependent tempera-387 ture correction, based on climatological temperature profiles, in the calculation of the

amfs[Bucsela et al., 2006; Boersma et al., 2002].

388

389 The lowest-elevation measurements have a large sensitivity to absorption in the bound-390 ary layer, while the zenith measurements are used as background reference spectra which 391 contain Fraunhofer structures and the stratospheric absorption features. Since photon 392 scattering largely occurs below the tropopause, the photons collected from different el-393 evation angles have essentially the same stratospheric path, but different light paths in 394 the troposphere. The difference between successive off-axis line-of-sight (los) and zenith 395 measurements is therefore only sensitive to the troposphere. For NO_{2} retrieval, radiative 396 transfer simulations show that under polluted conditions, the stratospheric contamination 397 is generally smaller than 1%. A more in-depth description of the max-doas measure-

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 23

398 ments, as they were done at the dandelions campaign, can be found in Brinksma et al.

[2007].

399

400 During the dandelions campaigns [Brinksma et al., 2007], various max-doas instru-401 ments operated quasi continuously from the Cabauw Experimental Site for Atmospheric 402 Research [Russchenberg et al., 2005] throughout May through mid July 2005 and through-403 out September 2006. These instruments were provided and operated by BIRA-IASB, the

University of Bremen, and the University of Heidelberg.

404

405 The Heidelberg max-doas instrument differs from the others in that it has a set of 406 three movable telescopes, which enable simultaneous measurement cycles in three azimuth

viewing directions Wagner et al. [2004]).

407

408 For the Bremen instrument [Wittrock et al., 2004], the zenith direction is viewed without 409 a mirror, while the other elevation angles in the measurement cycle are selected through a 410 rotating mirror inside the telescope. The range of elevation angles is 0^{∘} − 30^{∘}. In the 2006 411 dandelions campaign, separate uv and vis instruments were operated by the Bremen

group.

412

4.2.2. Agreement Between MAX-DOAS Instruments

413

414 The level of agreement achieved between the max-doas instruments is quantitatively 415 summarized in correlation plots (Fig. 4) where tropospheric NO_{2} columns from the BIRA 416 instrument are compared to those from the other groups, for the 2005 campaign. This 417 comparison is useful because there are differences between the instruments and their data 418 reduction procedures, and because instruments viewing different directions can be used 419 to detect directional dependences of the measurements. Very good agreement is found 420 between the BIRA and Bremen data sets (correlation coefficients of 0.9 and slope of 1.1),

X - 24 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

421 and also between BIRA and the three Heidelberg telescopes (correlations between 0.82 422 and 0.91), especially considering that the BIRA instrument was 200 m away from the 423 other two instruments. 424 When the BIRA and Heidelberg data sets were re-retrieved using identical NO_{2} absorp-425 tion cross-sections, an even better agreement is found, reaching a correlation coefficient 426 of 0.92 and a slope of 0.99. This level of agreement is only achieved when considering 427 the southwest-pointing Heidelberg telescope: this is approximately in the same direction 428 as the two other instruments. This highlights the degree to which horizontal inhomo-429 geneities in the NO_{2} field can strongly affect the agreement between ground based and 430 satellite based measurements.

4.2.3. Heterogeneity of the NO_{2} Field

431

432 If the tropospheric NO_{2} layer were horizontally homogenous, the observed NO_{2} scds for 433 the different azimuth angles observed from the three telescopes of the Heidelberg instru-434 ment would have been similar. The horizontal inhomogeneity of the NO_{2} concentration 435 field can be estimated from the observed differences among the various viewing directions. 436 Estimating the horizontal inhomogeneity is very important for the validation of satellite 437 instruments with ground based observations. In the presence of strong horizontal gradi-438 ents, ground based observations may not be representative for the average value within a

satellite ground pixel.

439

440 Since the horizontal extent of the absorption paths along the line of sight is largest 441 for low telescope elevation, those at 3^{∘} were used to estimate the heterogeneity of the 442 tropospheric NO_{2} concentration field. This was done by evaluating the scd in the three 443 azimuthal viewing directions at 3^{∘} elevation, and calculating the ratio of the maximum

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 25

444 and the minimum. A horizontally homogenous concentration field yields a ratio of one; the 445 more this ratio deviates from unity, the larger are the horizontal gradients. In addition to 446 the strength of the horizontal gradients, the direction of the NO_{2} gradient was estimated, 447 though in a limited way, since the Heidelberg max-doas was measuring in only three 448 azimuth directions. Fig. 5 displays the time series of the ratios at daily noon. High ratios 449 indicate strong gradients, and the color of the points indicates the direction of positive

gradient.

450

451 To interpret the retrieved information on the gradient of the tropospheric NO_{2} concen-452 tration field, it is important to consider two effects that can affect the observed scds, 453 especially for low elevation angles: First, the sensitivity to the relative azimuth angle 454 (between the telescope and the sun). This dependency becomes more pronounced for 455 increasing solar zenith angle (sza) and increasing aerosol load [Wagner et al., 2004]. Sec-456 ond, the effect of the atmospheric aerosol load on the atmospheric visibility, and thus on 457 the horizontal extents of the absorption paths along the line of sight. Thus, depending 458 on the aerosol load, the calculated ratio represents information on gradients over areas 459 of different horizontal extent. The dependence on the azimuth angle was found to be 460 below 15%, for sza between 20^{∘} and 80^{∘}. Almost all observed ratios of the maximum and 461 minimum NO_{2} scds (see Fig. 5) were much larger than this. Effective path lengths are 462 enhanced by aerosols above about 1 km, and diminished by aerosols below 1 km. For an 463 elevation angle of 3^{∘}, the effective path length is about 19 km in a pure Rayleigh-scattering 464 atmosphere, but this can be reduced to as little as 4.5 km by surface-level aerosols, or

X - 26 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

465 enhanced to 25 km by higher-altitude aerosols [Brinksma et al., 2007; Deutschmann and

Wagner, 2006; Wagner et al., 2007, 2004].

466

4.2.4. Comparisons with OMI tropospheric NO_{2}

467

468 The different max-doas data sets are compared to the omi Level-2 cloud-free data 469 (O_{2}−O_{2} cloud fractions in the omi products less than 20%) for 2005, in order to produce 470 the correlation plot presented in Figure 6 and Table 2. The max-doas data were linearly

interpolated to the satellite overpass time.

471

472 The omi tropospheric vertical columns were generally distributed from 0 to about 2.5× 473 10^{16} cm^{−2}. In one case the ground based max-doas column (3 × 10^{16} cm^{−2}) significantly 474 exceeded the corresponding satellite values, possibly due to a local enhancement of the 475 NO_{2} concentration at Cabauw. Because of its obvious anomaly, this point has been

excluded from the regression analysis.

476

477 The regression analyses show that similar results were achieved with the BIRA and 478 the Bremen data sets, the correlation coefficient between ground based and satellite data 479 being about 0.6. A lower correlation was obtained with the Heidelberg data when con-480 sidering only the southwest direction measurements (closest to the viewing direction of 481 both Bremen and BIRA instruments), possibly due to the smaller number of coincidences 482 with this instrument, and also the shorter integration time used, which may increase the 483 sensitivity to local inhomogeneities in the NO_{2} field. In order to further explore the im-484 pact of possible horizontal smoothing effects on the comparison results, the Heidelberg 485 measurements simultaneously recorded from all three directions have been averaged and 486 again compared with satellite data. The resulting correlation coefficients, also given in 487 Table 2, have significantly improved and are now the highest of the three max-doas in-

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 27

488 struments. This suggests that the scatter in max-doas versus satellite comparisons is,

indeed, largely dominated by the difference in their spatial and temporal averaging.

489

490 As is evident from the regression results, the omi tropospheric NO_{2} columns seem to be 491 systematically lower than the max-doas results, for both omi products considered. How-492 ever, it must be noted that the correlation coefficients are rather poor in all cases, which 493 might be due to several reasons including uncertainties in both ground based (geometrical 494 approximation) and satellite retrievals (amf sensitivity to errors in aerosols, clouds and 495 NO_{2} profile shape). However, the slopes are poorly determined, due to the small number

of poiints and the large data scatter.

496

4.3. Total column 497 4.3.1. Brewer 498 Cede et al. [2006] have described a method for retrieval of total-column NO_{2} from direct-499 sun measurements using a Brewer MK-III double monochromator spectrophotometer. The 500 Brewer MK-III instrument was primarily designed to make measurements of ozone from 501 wavelengths below 320 nm, and can measure spectral irradiance and radiance from 285 502 to 365 nm. Its measurement modes include a spectral scan mode, where the gratings 503 are moved and any wavelength can be selected, and a slit mask mode, in which a slit 504 mask is introduced in the optical path allowing nearly simultaneous measurements at 6 505 wavelengths, spaced about 3 nm apart. The spacing of the slits in the slit mask was 506 chosen to optimize the ozone retrievals between 303 and 320 nm, but in the 345–365 nm 507 range the measured wavelengths fall very nearly on maxima and minima in the NO_{2} 508 absorption spectrum (see Figure 7), which permits the retrieval of total column NO_{2}. 509 These measurements have been made at the NASA Goddard Space Flight Center, on a

X - 28 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

510 nearly continual basis since August 2004, with measurements made every half hour during 511 the sunlit hours. 512 The retrieved NO_{2} columns have a large instrumental noise, so data must be averaged 513 over several hours time in order to make meaningful comparisons to the omi-measured 514 values. However, the location of the instrument, 3 km from the Washington Capital 515 Beltway and 2 km from the Baltimore-Washington Parkway, on the outskirts of a major 516 metropolitan area, is such that there are often substantial sub-hour time variations in the 517 actual tropospheric NO_{2} concentrations. The combination of the intrinsic variability of the 518 measurements with the frequent occurrence of significant actual concentration variations 519 within a given time-window used for collocation with omi overpasses complicates the 520 process of using the Brewer data for validation of omi NO_{2} measurements. Comparisons 521 having useful statistical significance can be made using monthly averages of the Brewer 522 and omi datasets. 523 Figure 8 shows the comparison between the monthly mean Brewer-measured and omi-524 measured NO_{2} columns. In the Washington DC area, early afternoon NO_{2} columns are 525 dominated by the boundary layer columns. The difference that is seen, with omi-measured 526 columns that are about 35% smaller than the Brewer-measured columns, can thus be 527 largely attributed to the tropospheric NO_{2}. In Figure 9 the daily and monthly mean 528 values are plotted, along with the line of linear regression to the monthly means. The 529 regression analysis, performed on the monthly means, and weighted according to the

standard deviations, gives a slope of 0.67, with a correlation coefficient R = 0.95.

530

4.3.2. MultiFunction DOAS (MFDOAS) measurements

531

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 29

532 The mf-doas instrument observes scattered skylight with a 1^{∘} vertical fov at varying 533 viewing azimuth and elevation angles, as well as direct sunlight in the UV-visible spectral 534 region. From these measurements are retrieved NO_{2}, O_{3}, SO_{2}, and CH_{2}O slant columns. 535 The mf-doas spectrograph is a single pass commercial Czerny-Turner spectrograph of 536 focal length 300 mm. The instrument covers a wavelength range from 280 nm to 490 nm 537 with a spectral resolution of 0.82 nm (6 pixels fwhm). Scattered sky light is collected by 538 a 12 cm telescope and passes into the spectrograph through two filter wheels that contain 539 depolarizers, spectral flattening filters, and UV cutoff filters. Direct sunlight is fed into 540 a spectralon integrating sphere of diameter 8 cm before passing through the filter wheels 541 and results in a signal level similar to that from the scattered sky. A two-dimensional 542 ccd detector (512 × 2048 pixels) is used in the focal plane. Spectrograph stray light is 543 reduced by a spectral flattening filter, which reduces the long wavelength throughput of 544 the instrument relative to the short wavelength signal. A solar tracker moves the entire 545 instrument for positioning and sun tracking. An instrument schematic is presented in

Figure 10.

546

547 This ground-based mf-doas instrument was fielded in a prototype form during the 548 intex-b campaign for Aura/omi validation. It was positioned on the roof of a building at 549 Pacific Northwest National Laboratory in Richland, WA (pnnl; 46.3409^{∘}N, 119.2787^{∘}W), 550 located in an urban area known as Tri-Cities (the merged cities of Kennewick, Pasco and 551 Richland, WA) with total population of approximately 150,000 in an area of 250 km^{2}. 552 Pnnl is situated approximately 15 km north of the center of Richland, and northwest

X - 30 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

553 of the area’s population center. The major source of local NO_{2} pollution is vehicular 554 exhaust. 555 NO_{2} differential slant columns (dscd) were derived using the doas technique, based 556 on Beers law. A nonlinear least squares algorithm was used to fit our measured spectral 557 cross sections of NO_{2}, O_{3}, instrument spectral polarization, and Ring effect in the spectral 558 region 400–419 nm. A polynomial was included to remove slowly varying Rayleigh and 559 Mie scattering spectral shapes. The reference solar spectrum used for the data analysis 560 was measured at zenith at local noon on April 30, 2006, a day with very low pollution 561 levels. Raw spectra were corrected for detector dark background and flat field. Figure 562 11 shows typical residual optical densities after the least squares fitting procedure for 563 observations taken on May 9, 2006 at 5^{∘} elevation and 4 azimuth angles. The lidort 564 radiative transfer code [Spurr, 2001; Spurr et al., 2001] was used to calculate the air mass 565 factors (amf) to convert the dscd to vertical column density (vcd). As an example of 566 the results, Figure 12 presents the spatial and temporal variation of NO_{2} differential slant 567 column for May 9, a polluted day. Higher column densities were observed to the south and 568 east, toward the urban center, as expected. Measurements taken at 5^{∘} elevation showed 569 higher NO_{2} tropospheric column compared to 15^{∘} and 45^{∘} angles, as expected. These 570 elevated NO_{2} slant column densities were particularly pronounced during the morning 571 rush hour. 572 The Aura satellite flies over Tri-Cities area around 1330h with spatial resolution ap-573 proximately 13 km × 24 km. Figure 13 shows contour plots of omi tropospheric NO_{2} vcd 574 for May 9, derived from the Level-2 omi data product. omi tropospheric NO_{2} vertical 575 column densities “integrated” over several pixels in the mf-doas observation direction

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 31

576 were compared to mf-doas tropospheric NO_{2} vcd using a priori differential amfs for 577 clear days at pnnl from the lidort radiative transfer code. Figure 14 shows results for 578 the time period April 30 through May 13, 2006, with reasonable correlation observed for

these clear days.

579

580 The slope of the data in Figure 14 shows that omi determinations of tropospheric NO_{2} 581 vcd are 0.81 ± 0.11 of that determined from mf-doas with a correlation coefficient R^{2} of

0.92. Thus, omi measures a somewhat smaller vcd than that determined from mf-doas.

582

583 4.3.3. Pandora-1 Direct Sun DOAS measurements 584 The lightweight, portable Pandora-1 spectrometer system measures direct-sun irradi-585 ances from 270 to 500 nm at ∼ 0.5 nm resolution. The outdoor head sensor is mounted 586 on a tracking system and holds a single strand fiber optic cable, which collects the light 587 passed through a collimator (1.6^{∘} FWHM field of view) and a filter wheel. The other 588 end of the fiber is connected to a 75 mm focal length symmetric Czerny-Turner grating 589 spectrometer using a 1024 × 1 pixel cmos detector, stabilized to 20^{∘} ± 1. The total NO_{2} 590 column is retrieved by the doas method, in the 400–440 nm window, using a fixed ref-591 erence spectrum determined from Pandora-1 data obtained over an extended period of 592 at least 2 weeks. To estimate the NO_{2} amount in the reference spectrum, a bootstrap 593 method as described in Cede et al. [2006] was applied, on the assumption that a few 594 measurements were obtained when there were low tropospheric NO_{2} amounts (e.g., just 595 after sunrise). Figure 15 shows Pandora-1 data during the scout campaign in July 2006 596 at Thessaloniki, Greece. Excellent agreement is seen between the omi and Pandora-1 597 measurements, though the omi overpass times seem to occur just before or just after the 598 mid-day maximum in NO_{2} concentration; this limits the range of NO_{2} concentration val-

X - 32 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

599 ues explored in this comparison. A number of further field campaigns are planned, during

which Pandora-1/omi comparisons will be done.

600

4.3.4. Direct Sun DOAS (BIRA)

601

602 During the second dandelions campaign, a direct-sun doas instrument was operated 603 in addition to the max-doas instrument. The well-defined optical path and air mass 604 factor make this instrument equally sensitive to absorption along the whole optical path

and so provides accurate NO_{2} total columns.

605

606 The instrument is similar in concept to the max-doas: Inside the building, in a thermo 607 regulated box, a grating spectrometer covering the UV-Vis region is coupled to a cooled 608 ccd detector, connected by depolarizing fiber optic bundle to the external optical head. 609 Outside, alongside the max-doas scanning telescope, a collimating optic tube is mounted

on a brusag commercial sun-tracking system, holding the fiber.

610

611 The retrieval is also done using the doas approach: The ratios of the measured radiance 612 spectra to a reference spectrum are analyzed with respect to a set of reference spectra, 613 in the 425–450 nm window, including laboratory spectra of O_{3}, H_{2}O, O_{2}−O_{2}, the com-614 puted the Ring effect spectrum, and NO_{2} cross-sections at two different temperatures. In 615 contrast to the analysis of max-doas data, a fixed reference spectrum (measured on 7 616 September 2006) has been used for the whole time-series. The NO_{2} residual slant column 617 amount included in this reference spectrum has been obtained by analysing it with re-618 spect to the Kurucz solar atlas [Kurucz et al., 1984], which was assumed to be free of NO_{2} 619 absorption. The Kurucz solar spectrum was convolved with a precisely measured instru-620 ment slit function to match the instrument’s spectral resolution. Based on this analysis,

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 33

621 total absolute slant columns could be derived from direct sun measurements; these were

transformed into total vertical columns using geometrical amfs.

622

623 Figure 16 shows the time series for the bira doas measurements of the total column 624 NO_{2} (filled dots), which provides a good idea of the diurnal variation of NO_{2} levels. 625 The open squares show the collocated omi measurements (one or two per day). The 626 omi snapshots of vertical column NO_{2}, for the most part, appear to be in quite good 627 agreement with the ground-based measurements. Note that the omi data are filtered for

clouds (cloud fraction ≤ 20%).

628

629 Figure 17 shows the correlation plot of the collocated data (the point nearest in time 630 to the omi overpass). A linear regression, constrained to pass through the origin, gives a 631 slope of 0.84 ± 0.05. The scatter in the data (R = 0.68), and the relatively small number 632 of data points (N = 26)do not permit a statistically significant estimation of an additive 633 bias.

4.3.5. FTUVS measurements at Table Mountain, California

634

635 Another instrument that has been used to validate omi NO_{2} total column measurements 636 uses the Fourier Transform Ultraviolet Visible Spectrometer (ftuvs), a uv-vis-nir in-637 terferometer, at the Table Mountain Facility (tmf), north of Los Angeles, California, at 638 34^{∘}22.9′ N, 117^{∘}40.8′ W, at an altitude of 2290 m (7300′) [Cageao et al., 2001]. Spectra are 639 recorded in the direct solar absorption mode with a spectral resolution of 0.0013 nm, which 640 is sufficient to resolve NO_{2} vibronic features. By measuring the doppler-shifted spectra 641 from the east and west solar limbs, and taking the ratio of the two, one can remove the

X - 34 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

642 solar Fraunhofer lines; there is no need to measure a high-sun reference spectrum, as in a 643 number of the other methods described in this overview. 644 The instrument is not readily transportable. The ftuvs observation site overlooks the 645 Antelope Valley, north of the Los Angeles Basin. This area is characterized by relatively 646 clean air under most conditions, but is often influenced by polluted air from Los Angeles 647 in the afternoon, advected through the Cajon Pass. While considerably above the tro-648 pospheric background under these conditions, the NO_{2} column abundance values rarely 649 exceed 1 × 10^{16}molecules cm^{−2}, which is considerably smaller than values measured di-650 rectly downwind of a polluted urban area (see Fig. 8). Because the altitude of tmf is 651 about 2500 feet above the Antelope Valley, ftuvs column abundance measurements of 652 NO_{2} will be biased relative to the center of the omi footprint. The bias is small relative 653 to the total column, and will not have a significant effect on the slope of the omi-ftuvs 654 correlation. The omi data used for validation were sorted by distance from the tmf 655 site, in order to mitigate somewhat the possible effects of the distribution of elevations 656 within a fov. It was found that a minimum distance of about 10 km is required for good

intercomparison.

657

658 The ftuvs instrument time is shared with other Aura validation activities. On average,

measurements were acquired twice a week over the period March–November, 2006.

659

660 The slant column NO_{2} amounts are retrieved by fitting the measured absorption spectra 661 to laboratory spectra at a number of temperatures [Nizkorodov et al., 2004], in windows 662 containing 10 to 20 NO_{2} rotational lines. Geometric amfs were used to convert the 663 scds to vcds. Figure 18 presents the comparison of the omi-derived and ftuvs-derived 664 measurements of total column NO_{2}. In this figure, the points where the omi fov center

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 35

665 fell within 10 km of the Table Mountain Facility site are colored red. The linear regression 666 line shown is fit only to those points. As shown, this line has a slope of 0.77 ± 0.41, and 667 it does not go through the origin. This data set suggests that the omi NO_{2} totals are 668 underestimated in the middle of the data range, but that there may also be a positive 669 additive bias.

4.4. NO_{2} Profile measurements 670 As pointed out in previous sections, and in Boersma et al. [2002] and Bucsela et al. 671 [2006], the shape of the vertical profile of NO_{2} influences the (physical) air mass fac-672 tors. The omi NO_{2} algorithm uses a set of assumed profiles, which were derived from 673 model studies; these assumed profiles thus affect the retrieved total and tropospheric NO_{2} 674 amounts. It is therefore important to evaluate how well the assumed profiles approximate 675 the actual profiles, vis-à-vis the air mass factor calculation. There have been very few 676 efforts to measure NO_{2} profiles [Heland et al., 2002; Martin et al., 2006]. Recent efforts 677 include measurements during the September 2006 dandelions campaign (lidar, in situ 678 at two altitudes, and max-doas at two altitudes, see section 4.4.1), and aircraft-based in

situ measurements taken during the INTEX-B campaign in North America.

679

4.4.1. NO_{2} lidar 680

681 NO_{2} profiles were measured by a lidar system, developed at rivm, during the dan-682 delions campaign in September 2006 [Brinksma et al., 2007]. The lidar consists of an 683 emitter and a receiver unit. The entire system is housed in a truck, constituting a fully 684 self-supporting mobile laboratory. The emitter unit consists of a pulsed pump laser-dye 685 laser combination, running at 30 Hz. The dye laser is tuned to 449.10 nm and detuned 686 to 448.31 nm every other pulse. The latter wavelength is absorbed more strongly by NO_{2}

X - 36 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

687 than the former. The laser pulses, 40 mJ in energy, 10 ns in duration, are directed into the 688 atmosphere, where they are scattered by gas molecules and aerosol particles. The receiver 689 unit collects the backscattered light, through a 280 mm telescope, onto a photomultiplier 690 tube, with an interference filter to block daylight. A digitizer samples the signals with a

range resolution of 3.75 m.

691

692 The NO_{2} concentration at a certain altitude is derived from the log of the ratio of 693 the backscattered signals at the two wavelengths, using the differential absorption lidar 694 (dial) method. Since the laser pulses are not emitted from the center of the telescope, the 695 laser beam is not in view of the telescope at close range, and thus the lidar is effectively 696 blind for the first 500 m. When measurements starting near the surface are required, the 697 emitter section and receiving telescope are tilted through various elevation angles; the 698 measurements are combined into a single profile, where elevations close to the horizontal 699 yield NO_{2} concentrations at low altitudes but pertaining to a certain horizontal extent 700 away from the instrument (for a near-horizontal measurement, typically about 2500 m), 701 while a zenith observation is performed exactly above the truck. Completing one vertical 702 profile typically takes 50 minutes, providing data in a altitude range of a few meters up to 703 approximately 2500 m, with an accuracy of 0.2−0.4 μg m^{−3}. Range and accuracy depend 704 on atmospheric conditions. The vertical resolution of a profile varies, and typically is 705 about 15 m at the lowest altitude, increasing to over 500 m at the highest altitude. The

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 37

706 resolution arises from averaging of data over an altitude range, based on signal-to-noise 707 considerations. 708 A paper describing the lidar and other time-resolved three dimensional observations 709 of NO_{2} during the 2006 dandelions campaign is in preparation (H. Volten et al., in

preparation).

710

711 Figure 19 presents examples of profile measurements for a relatively clean day, Septem-712 ber 9, 2006, and for a polluted day, September 12, 2006; in both cases, there was little-713 to-no cloud cover. The concentration of NO_{2} is high at ground level, and drops to zero 714 (within the accuracy of the measurement) above the boundary layer. The boundary layer 715 heights, provided by the boundary layer lidar at Cabauw, are indicated in Figure 19 by a 716 dashed line. The figure shows that the day-to-day variations in NO_{2} at the surface may 717 be considerable, from around 3 μg NO_{2} m^{−3} on a clean day to more than 50 μg NO_{2} m^{−3}

on a polluted day. Large diurnal variations may also occur.

718

719 4.4.2. In-situ aircraft measurements 720 In situ measurements of NO_{2} from the DC-8 aircraft were obtained during the intex-a 721 (summer 2004), pave (winter 2005) and intex-b (spring 2006) campaigns. These have 722 been discussed by Bucsela et al. [2007]. The NO_{2} profiles from these experiments are useful 723 for validating both the shapes of the model profiles used in the omi retrieval algorithm, 724 and, in turn, the tropospheric column amounts from the satellite retrievals. The aircraft 725 profiles obtained during intex-a and pave were combined into composite land and ocean 726 profiles. The in situ profiles were seen to be very similar to the annual mean geos-chem 727 profiles used to retrive tropospheric NO_{2} columns from omi, and the amfs computed from 728 the measured profiles were slightly larger than those calculated using the model profiles.

X - 38 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

729 A more quantitative analysis was performed using a set of approximately 70 profiles 730 measured during intex-b. Error-weighted linear regressions comparing the amfs yielded 731 a slope of 1.10 ± 0.10 (in situ profile amf greater than that used by the omi algorithm).

This means that the omi vcd would overestimate the actual vcd by 10%(±10%).

732

733 In situ measurements of NO_{2} can also be used to validate tropospheric column amounts 734 from omi. The intex-b data were used for this analysis by Bucsela et al. [2007] (see 735 also Boersma et al. [2007]). Two representative profile analyses are shown in Figure 20. 736 The full set of profiles from intex-b where used The correlation between the aircraft 737 and omi data sets was good (R = 0.83). This comparison is shown in Figure 21. The 738 integrated in situ tropospheric columns were found to be somewhat larger than the omi 739 Level-2 columns, as indicated by the slope of 1.10 ± 0.08. Although some of the in situ 740 columns required significant extrapolations, sensitivity studies indicated that the overall 741 results were generally robust with respect to the choices made for the profile binning, 742 integration and extrapolation, as well as being relatively insensitive to the errors assumed 743 for the weights. The insensitivity to extrapolation is consistent with findings in a similar

aircraft study by Heland et al. [2002].

744

5. Conclusions and discussion 745 This paper has presented a number of results of experiments where ground- and aircraft-746 based measurements of NO_{2} can be compared with collocated measurements and retrieval 747 by omi. Since some measurements estimate the stratospheric column, others the tropo-748 spheric column, and still others the total column, their results can be used to validate the

omi NO_{2} standard data product’s estimates of these columns.

749

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 39

750 Table 3 summarizes the results of the numerous validation studies that have been dis-751 cussed in this overview. 752 On the basis of the saoz and doas measurements, which are most sensitive to the 753 stratospheric NO_{2} columns, the omi stratospheric NO_{2} appears to agree with the ground-754 based measurements to within ∼ 10%. 755 The omi tropospheric column appears to be consistently lower than the various ground-756 based measurements, though there is some inconsistency amongst those ground-based 757 measurements. Though many of the various instruments and methods for measuring 758 tropospheric and total NO_{2} have not themselves been validated, it is noteworthy that they 759 all give NO_{2} estimates that are on average greater than those retrieved from omi. This 760 may indicate a bias in the omi retrieval. However, a number of cases have been studied, 761 where average differences between omi and ground based measurements decrease as the 762 geographic match up criterion is tightened. It is likely to be due to the inhomogeneity of 763 the tropospheric NO_{2} field, and, in particular, the fact that ground-based measurements 764 are often made in or near regions of moderate to strong sources of NO_{2}: The omi fov 765 that includes the site will also include a substantial (∼ 10^{2} km^{2}) regions where much lower 766 NO_{2} concentrations prevail. The was borne out in the Brewer studies [Cede et al., 2006] 767 and in the correlation studies of Veefkind et al. [2007] (see also Section 1.) However, the 768 data taken at tmf (Section 4.3.5) were mostly obtained under conditions of relatively

clean tropospheric air, and these data also suggested a negative bias for omi retrieval.

769

770 Potential biases can arise at any of the steps in the algorithm. Instrumental artifacts 771 are known to give rise to the cross-track bias (striping) and the destriping process can 772 certainly give rise to a general bias. The stratosphere-troposphere separation is based

X - 40 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

773 on an initial amf, and any bias in that amf will result in a bias in the background 774 (mostly stratospheric) field. After the stratosphere-troposphere separation, a new amf 775 is constructed, based on model-based-climatology derived a priori profiles. The aircraft 776 in situ measurements of NO_{2} profile shape suggest that the a priori profile shapes are 777 essentially correct, in that the two do not give appreciably different amfs. The amf 778 is also sensitive to the surface albedo. The omi algorithm uses a climatological surface

albedo, and this may be a worthy subject for future validation studies.

779

780 It should be mentioned that all the validation studies reviewed here focused on mostly 781 cloud-free conditions. However, while omi fovs are considerably smaller than those of 782 earlier atmospheric remote sensing instruments, they are still large enough that very few

can be expected to be completely uncontaminated by clouds.

783

Acknowledgments

784

785 Part of this research was carried out at the Jet Propulsion Laboratory, California Insti-786 tute of Technology, under contract with the National Aeronautics and Space Administra-787 tion.

References

References 788 Bey, I., D. J. Jacob, R. M. Yantosca, J. A. Logan, B. D. Field, A. M. Fiore, Q. Li, H. Y. 789 Liu, L. J. Mickley, and M. G. Schultz (2001), Global modeling of tropospheric chem-790 istry with assimilated meteorology: Model description and evaluation, J. Geophys.

Res., 106, 23 073–23 096.

791

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 41

792 Boersma, K. F., E. Bucsela, E. Brinksma, and J. F. Gleason (2002), NO_{2}, in OMI 793 Algorithm Theoretical Basis Document, Volume IV: Trace Gas Algorithms, edited by

K. Chance, KNMI, NASA/GSFC, FMI, pp. 15–36.

794

795 Boersma, K. F., H. J. Eskes, and E. J. Brinksma (2004), Error analysis for tropospheric

NO_{2} retrieval from space, J. Geophys. Res., 109, doi:10.1029/2003JD003962, D04 311.

796

797 Boersma, K. F., D. J. Jacob, E. J. Bucsela, A. E. Perring, R. Dirksen, R. J. van der A, 798 R. M. Yantosca, R. J. Park, and R. C. Cohen (2007), Validation of OMI tropospheric 799 NO_{2} observations during INTEX-B and application to constrain NOx emissions over

the eastern United States and Mexico, J. Geophys. Res., —, doi:---, —.

800

801 Bovensmann, H., J. P. Burrows, M. Buchwitz, J. Frerick, V. V. Rozanov, K. V. Chance, 802 and A. P. H. Goede (1999), SCIAMACHY: Mission objectives and measurement

modes, J. Atmos. Sci., 56, 127–150.

803

804 Brinksma, E. J., G. Pinardi, R. Braak, H. Volten, A. Richter, A. Schönhardt, M. van 805 Roozendael, C. Fayt, C. Hermans, R. J. Dirksen, T. Vlemmix, A. J. C. Berkhout, 806 D. P. J. Swart, H.Oetjen, F. Wittrock, T. Wagner, O. W. Ibrahim, G. de Leeuw, 807 M. Moerman, R. L. Curier, E. A. Celarier, W. H. Knap, J. P. Veefkind, H. J. Eskes, 808 M. Allaart, R. Rothe, A. J. M. Piters, and P. F. Levelt (2007), The 2005 and 2006

DANDELIONS NO_{2} and aerosol validation campaigns, J. Geophys. Res., —, —.

809

810 Bucsela, E., A. E. Perring, R. C. Cohen, J. F. Gleason, K. F. Boersma, M. O. Wenig, 811 T. H. Bertram, P. J. Wooldrige, R. Dirksen, E. A. Celarier, and J. Veefkind (2007), 812 A comparison of NO_{2} in situ aircraft measurements with data from the Ozone Mon-

itoring Instrument, J. Geophys. Res., —, doi:---, —.

813

X - 42 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

814 Bucsela, E. J., E. A. Celarier, M. O. Wenig, J. F. Gleason, J. P. Veefkind, K. F. 815 Booersma, and E. J. Brinksma (2006), Algorithm for NO_{2} vertical column retrieval 816 from the Ozone Monitoring Instrument, IEEE Trans. Geosci. Rem. Sens., 44 (5),

doi:10.1109/TGRS.2005.863715, 1245–1258.

817

818 Burrows, J. P., A. Richter, A. Dehn, B. Deters, S. Himmelmann, S. Voigt, and J. Orphal 819 (1999a), Atmospheric remote sensing reference data from GOME-2. temperature-820 dependent absorption cross-sections of O_{3} in the 231-794 nm range, J. Quant. Spec-

trosc. Radiat. Transfer, 61, 509517.

821

822 Burrows, J. P., M. Weber, M. Buchwitz, V. V. Rozanov, A. Ladstatter-Weissenmayer, 823 A. Richter, R. DeBeek, R. Hoogen, K. Bramstedt, and K. U. Eichmann (1999b), The 824 global ozone monitoring experiment (GOME): Mission concept and first scientific

results, J. Atmos. Sci., 56, 151–175.

825

826 Cageao, R. P., J. F. Blavier, J. P. McGuire, Y. B. Jiang, V. Nemtchinov, F. P. Mills, and 827 S. P. Sander (2001), High resolution fourier transform ultraviolet-visible spectrometer 828 for the measurement of atmospheric trace species: Application to OH, Appl. Opt., 40, 829 2024–2030. 830 Cede, A., J. Herman, R. Richter, N. Krotkov, and J. Burrows (2006), Measurements of 831 nitrogen dioxide total column amounts using a brewer double spectrophotometer in

direct sun mode, J. Geophys. Res., 111, doi:10.1029/2005JD006585.

832

833 Celarier, E. A., J. F. Gleason, E. J. Bucsela, K. F. Boersma, E. Brinksma, J. P. Veefkind, 834 and P. Levelt (2006), OMNO2 README file, Tech. rep., NASA Goddard Space Flight

      Center. URL http://toms.gsfc.nasa.gov/omi/no2/OMNO2_readme.pdf.

835

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 43

836 Chance, K. V. and R. J. D. Spurr (1997), Ring effect studies: Rayleigh scattering, 837 including molecular parameters for rotational Raman scattering, and the Fraunhofer

spectrum, Appl. Opt., 36 (21), 5224–5230.

838

839 Cleary, P. A., P. J. Wooldridge, and R. C. Cohen (2002), Laser-induced fluorescence de-840 tection of atmospheric NO_{2} with a commercial diode laser and a supersonic expansion,

Appl. Opt., 41, 6950–6956.

841

842 Denis, L. e. a. (2005), A new software suite for NO_{2} vertical profile retrieval from 843 ground-based zenith-sky spectrometers, J. Quant. Spectrosc. Radiat. Transfer, 92 (3), 844 321–333. 845 Deutschmann, T. and T. Wagner (2006), TRACY-II Users manual, University of Hei-

delberg.

846

847 Dobber, M., R. Dirksen, P. Levelt, G. van den Oord, R. Voors, Q. Kleipool, G. Jaross, 848 M. Kowalewski, E. Hilsenrath, G. Leppelmeier, J. de Vries, W. Dierssen, and N. Roze-849 meijer (2006), Ozone monitoring instrument calibration, IEEE Trans. Geosci. Rem.

Sens., 44 (5), doi:10.1109/TGRS.2006.869987, 1209–1238.

850

851 Finlayson-Pitts, B. J. and J. Pitts, James N. (2000), Chemistry of the Upper and Lower 852 Atmosphere: Theory, Experiments, and Applications, Academic Press, ISBN 012-853 257060X. 854 Gordley, L., J. Russell, L. Mickley, J. Frederick, J. Park, K. Stone, G. Beaver, J. McIn-855 erney, L. Deaver, G. Toon, F. Murcray, R. Blatherwick, M. Gunson, J. Abbatt, 856 R. Mauldin, G. Mount, B. Sen, and J. Blavier (1996), Validation of nitric oxide 857 and nitrogen dioxide measurements made by the halogen occultation experiment for

UARS platform, J. Geophys. Res., 101 (D6), doi:A1996UJ40400034, 10 241–10 266.

858

X - 44 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

859 Heland, J., H. Schlager, A. Richter, and J. P. Burrows (2002), First comparison of 860 tropospheric NO_{2} column densities retrieved from GOME measurements and in situ

aircraft profile measurements, Geophys. Res. Lett., 29 (20), Article No. 1983.

861

862 Ionov, D., F. Goutail, and J. Pommereau (2006a), Validation of satellite data on total 863 NO_{2}: GOME, SCIAMACHY and OMI nadir viewing instruments compared to UV-864 visible SAOZ network, in Proc. of 3rd International DOAS Workshop, Bremen, 20-22

March, 2006, pp. —.

865

866 Ionov, D., F. Goutail, J.-P. Pommereau, A. Bazureau, E. Kyro, T. Portafaix, G. Held, 867 P. Ericksen, and V. Dorokhov (2006b), Ten years of NO_{2} comparisons between ground-868 based SAOZ and satellite instruments (GOME, SCIAMACHY, OMI), in Atmospheric

Science Conference, ESRIN, Frascati, Italy, 8-12 May 2006, ESA SP-628, pp. —.

869

870 Ionov, D., Y. Timofeyev, F. Goutail, J.-P. Pommereau, and A. Shalamyansky (2007), 871 Delta-validation of ENVISAT SCIAMACHY total ozone and NO_{2} with the data of 872 ground-based UV-VIS measurements (m-124 and SAOZ), in 3rd Workshop on the 873 Atmospheric Chemistry Validation of Envisat (ACVE-3), ESRIN, Italy, 4-7 December

2006, ESA SP-642, pp. —.

874

875 Kurucz, R. L., I. Furenlid, J. Brault, and L. Testerman (1984), Solar flux atlas from 876 296 nm to 1300 nm, vol. National Solar Observatory Atlas No. 1, National Solar

Observatory.

877

878 Lambert, J., M. Van Roozendael, P. Simon, J. Pommereau, F. Goutail, J. Glea-879 son, S. Andersen, D. Arlander, N. Van, H. Claude, J. de la Noe, M. De Maziere, 880 V. Dorokhov, P. Eriksen, A. Green, K. Tornkvist, B. Hoiskar, E. Kyro, J. Lev-881 eau, M. Merienne, G. Milinevsky, H. Roscoe, A. Sarkissian, J. Shanklin, J. Stahe-

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 45

882 lin, C. Tellefsen, and G. Vaughan (2001), Combined characterisation of GOME and 883 TOMS total ozone measurements from space using ground-based observations from

the NDSC, Adv. Space Res., 26 (12), 1931–1940.

884

885 Levelt, P. F. and P. K. Bhartia (2007), Tbd, J. Geophys. Res., TBD(TBD), doi:TBD, 886 TBD. 887 Levelt, P. F., E. Hilsenrath, G. W. Leppelmeier, G. H. J. van den Oord, P. K. Bhartia, 888 J. Taminnen, J. F. de Haan, and J. P. Veefkind (2006a), Science objectives of the 889 Ozone Monitoring Instrument, IEEE Trans. Geosci. Rem. Sens., 44 (5), doi:10.1109/

TGRS.2006.872336, 1199–1208.

890

891 Levelt, P. F., G. H. J. van den Oord, M. R. Dobber, A. Malkki, H. Visser, J. de Vries, 892 P. Stammes, J. Lundell, and H. Saari (2006b), The Ozone Monitoring Instrument,

IEEE Trans. Geosci. Rem. Sens., 44 (5), doi:10.1109/TGRS.2006.872333, 1093–1101.

893

894 Martin, R., C. Sioris, K. Chance, T. Ryerson, T. Bertram, P. Woolridge, R. Cohen, 895 J. Neuman, A. Swanson, and F. Flocke (2006), Evaluation of space-based constraints 896 on nitrogen oxide emissions with regional aircraft measurements over and downwind

of eastern north america, J. Geophys. Res., 111, doi:10.1029/2005JD006680.

897

898 Martin, R. V., D. J. Jacob, J. A. Logan, I. Bey, R. M. Yantosca, A. C. Staudt, Q. Li, 899 A. M. Fiore, B. N. Duncan, H. Liu, P. Ginoux, and V. Thouret (2002), Interpretation 900 of TOMS observations of tropical tropospheric ozone with a global model and in situ

observations, J. Geophys. Res., 107 (D18), Article No. 4351.

901

902 Nizkorodov, S. A., S. P. Sander, and L. R. Brown (2004), Temperature and pressure 903 dependence of high-resolution air-broadened absorption cross sections of NO_{2} (415–

525 nm), J. Phys. Chem., A108, 4864–4872.

904

X - 46 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

905 Piters, A., K. Bramstedt, J. Lambert, and B. Kirchhoff (2006), Overview of sciamachy

validation: 2002-2004, Atmos. Chem. & Phys., 6, 127–148.

906

907 Platt, U. (1994), Chemical Analysis Series, 127, Differential optical absorption spec-908 troscopy (DOAS), Air monitoring by spectroscopic techniques, John Wiley & Sons, 909 Inc. 910 Russchenberg, H., F. Bosveld, D. Swart, H. ten Brink, G. de Leeuw, R. Uijlenhoet, 911 B. Arbesser-Rastburg, H. van der Marel, L. Ligthart, R. Boers, and A. Apituley 912 (2005), Groundbased atmospheric remote sensing in The Netherlands; European out-913 look, IEICE Transactions on Communications, E88-B(6), doi:10.1093/ietcom/e88-b.

6.2252, 2252–2258.

914

915 Spurr, R. J. D. (2001), Linearized radiative transfer theory: A general discrete ordi-916 nate approach to the calculation of radiances and analytic weighting functions, with 917 application to atmospheric remote sensing, Ph.D. thesis, Technical University of Eind-

hoven, The Netherlands.

918

919 Spurr, R. J. D., T. P. Kurosu, and K. V. Chance (2001), A linearized discrete ordinate 920 radiative transfer model for atmospheric remote sensing retrieval, J. Quant. Spectrosc.

Radiat. Transfer, 68, 689–735.

921

922 Thornton, J. A., P. J. Wooldridge, and R. C. Cohen (2000), Atmospheric NO_{2}: In 923 situ laser-induced fluorescence detection at parts per trillion mixing ratios, Analytical

Chemistry, 72, 528–539.

924

925 Vandaele, A. C., C. Hermans, P. C. Simon, M. Carleer, R. Colin, S. Fally, M. F. 926 Mérienne, A. Jenouvrier, and B. Coquart (1998), Measurements of the NO_{2} absorp-927 tion cross-section from 42000 cm^{−1} to 10000 cm^{−1} (238–1000 nm) at 220 K and 294

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 47

K, J. Quant. Spectrosc. Radiat. Transfer, 59, 171–184.

928

929 Veefkind, J. P. and E. A. Celarier (2006), OMI level 2 NO2 data product specifica-930 tion, Tech. Rep. SD-OMIE-KNMI-352, KNMI and NASA Goddard Space Flight

931  Center. URL http://disc.sci.gsfc.nasa.gov/Aura/OMI/OMNO2_data_product_
      specification.pdf.

932

933 Veefkind, J. P., E. J. Brinksma, K. F. Boersma, H. Eskes, J. F. Gleason, E. J. Bucsela, 934 E. A. Celarier, M. O. Wenig, and P. F. Levelt (2007), High spatial resolution NO_{2} 935 observations over europe using data from the ozone monitoring instrument, Geophys.

Res. Lett. (To appear.).

936

937 Wagner, T., B. Dix, C. v. Friedeburg, U. Frieß, S. Sanghavi, R. Sinreich, and U. Platt 938 (2004), MAX-DOAS O4 measurements - a new technique to derive information on 939 atmospheric aerosols. (I) Principles and information content, J. Geophys. Res., 109,

doi:10.1029/2004JD004904.

940

941 Wagner, T., J. Burrows, T. Deutschmann, B. Dix, F. Hendrick, C. von Friedeburg, 942 U. Frieß, K. Heue, H. Irie, H. Iwabuchi, Y. Kanaya, J. Keller, C. A. McLinden, 943 H. Oetjen, E. Palazzi, A. Petritoli, U. Platt, O. Postylyakov, J. Pukite, A. Richter, 944 M. van Roozendael, A. Rozanov, V. Rozanov, R. Sinreich, S. Sanghavi, and F. Wit-945 trock (2007), Comparison of Box-Air-Mass-Factors and Radiances for Multiple-Axis 946 Differential Optical Absorption Spectroscopy (MAX-DOAS) Geometries calculated

from different UV/visible Radiative Transfer Models, Atm. Chem. Phys, accepted.

947

948 Wenig, M. O., A. M. Cede, E. J. Bucsela, E. A. Celarier, K. F. Boersma, J. P. Veefkind, 949 E. Brinksma, J. F. Gleason, and J. R. Herman (2007), Validation of omi tropospheric 950 no_{2} column densities using direct-sun mode brewer measurements at nasa goddard

X - 48 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

space flight center, J. Geophys. Res., —, doi:---, —.

951

952 Wittrock, F., H. Oetjen, A. Richter, S. Fietkau, T. Medeke, A. Rozanov, and J. P. 953 Burrows (2004), MAX-DOAS measurements of atmospheric trace gases in Ny-Ålesund

- Radiative transfer studies and their application, Atmos. Chem. Phys., 4, 955–966.

954

Figure 1. Geographical distribution of eight CNRS-operated saoz stations in the NDACC network. The colored field presented is the annual mean tropospheric NO_{2} column amounts, as measured by omi.

Figure 2. Simulated time-history of stratospheric NO_{2} at a mid-latitude station (OHP), and a high-latitude station (Scoresbysund), for spring and fall.

X - 50 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

Figure 3. Time series of the difference between omi and saoz-measured stratospheric NO_{2} in units of 10^{15}cm^{−2}. The sites are ordered from North to South.

D R A F T                   August 27, 2007, 12:40pm                   D R A F T

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 51

Table 1. Absolute average and r.m.s. difference (Δ, ρ; ×10^{15} cm^{−2}), and correlation (R) between ground-based saoz and satellite data, adjusted to sunrise omi total and stratospheric NO_{2} (2004-2006)

Station  OMI–SAOZ  OMI–SAOZ
total column  strat. column
Δ ρ R Δ ρ R
Scoresby Sund +0.02 0.82 0.93 −0.33 0.44 0.99
Sodankyla  +0.58 2.03 0.71 −0.28 0.56 0.97
Salekhard  +1.05 1.57 0.86 +0.24 0.73 0.95
OHP  +1.48 2.39 0.49 −0.71 0.90 0.74
Reunion  +0.17 0.73 0.31 −0.45 0.56 0.67
Bauru  +0.63 2.03 0.17 −0.80 0.97 0.58
Kerguelen  −0.04 0.58 0.88 −0.37 0.59 0.89
Dumont dUrville +0.27 1.10 0.88 −0.23 0.66 0.96
OVERALL:  +0.53 1.56 0.70 −0.37 0.70 0.92

X - 52 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

Figure 4. Scatter plots of the tropospheric NO_{2} columns retrieved during the 2005 campaign from the BIRA max-doas instrument and, respectively, the Bremen maxdoas (top left), the Heidelberg max-doas for the 3 pointing directions North-West (top right), South-East (bottom left) and South-West (bottom right). The regression analysis parameters are given in the legends. It has to be noted that agreement with the Heidelberg observations can be further improved after homogenisation of the retrieval settings (see text).

Figure 5. Maximum and minimum NO_{2} scd observed for an elevation angle of 3^{∘} of the Heidelberg max-doas telescopes observing under three different azimuth angles at Cabauw during the dandelions campaign in 2005. High ratios indicate large horizontal gradients of the tropospheric NO_{2} concentration field, colors indicate the direction of the gradient (directed toward higher values).

Figure 6. Correlations between tropospheric NO_{2} from the three max-doas instruments at Cabauw (BIRA, Bremen, and Heidelberg SW direction) and omi Level-2 OMI data are included if cloud fractions were less than 20%. Correlation and regression coefficients are summarized in Table 2. The dashed lines are provided as a visual aid, and show slopes of 1/2, 1, and 2. The data point in brackets is clearly an outlier (see text), and was not included in the regression analysis presented in table 2.

Table 2. Summary of statistical analyses of comparisons between tropospheric NO_{2} from max-doas data and omi Level-2.

N R Intercept Slope RMS diff. Relative RMS diff.
BIRA south  29 0.60 4.29  0.52  4.82  52%
Bremen southwest  29 0.63 3.93  0.59  4.44  48%
Heidelberg southwest  21 0.45 4.27  0.8  5.38  56%
Heidelberg spatial average 21 0.65 1.99  0.85  3.89  40%

N denotes number of collocations, and R is Pearson correlation coefficient. The intercept and (absolute) RMS difference are in units of 10^{15} m^{−2} Figure 7. Typical optical depths of the main trace gases in the Brewer MK-III wavelength range. Lower: NO2 optical depth for 1 DU (= 2.7 × 10^{16} cm^{−2}), 6 slit positions with noise estimates.

Figure 8. Comparison between the monthly mean Brewer-measured and omi-measured NO_{2} total columns.

Figure 9. Daily mean and monthly mean values of NO_{2} total column measured by the Brewer instrument and omi. The line of linear regression is also shown.

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 57

Figure 10. Schematic drawing of the mf-doas instrument. Figure 11. Representative residual optical densities for observations of May 9, 2006 at 5^{∘} elevation and 4 azimuth angles.

D R A F T                   August 27, 2007, 12:40pm                   D R A F T

Figure 12. Example of measured spatial and temporal variation of mf-doas-measured NO_{2} differential scd for May 9, 2006

Figure 13. Tropospheric NO_{2} vcd over the Tri-Cities area of Washington State on May 9, 2006. The contour map is derived from the individual omi fov measurements.

CELARIER ET AL.: VALIDATION OF OMI NO_{2} X - 59

Figure 14. Correlation between omi tropospheric NO_{2} vertical column and mf-doas NO_{2} tropospheric vcd for measurements made from April 30 through May 13, 2006

X - 60 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

Figure 15. Pandora-1 Data measured during the period 14–24 July 2006, in Thessaloniki (grey dots) and omi overpass data for the site (large black dots).

D R A F T                   August 27, 2007, 12:40pm                   D R A F T

Figure 16. Time series of NO_{2} vcd measured by the BIRA direct-sun doas instrument. The open squares show the collocated omi measurements.

Figure 17. Correlation plot showing the collocated omi and BIRA direct-sun doas instrument measurements (open squares in Figure 16).

Figure 18. omi versus ftuvs measurements of total column NO_{2}, binned by distance between TMF and the centroid of the omi fov (distance indicated by color, see inset scale)

Figure 19. Lidar NO_{2} profile (black circles) and NO_{2} monitor value (open circle) measured at Cabauw. Horizontal bars indicate two-sigma values for the concentration. For the lidar data, vertical bars indicate the height intervals over which concentrations have been determined. The boundary layer height is indicated by a dashed line. Top panel: Clean day, September 9, 2006 (lidar measurement from 12:04–12:52 UT). On this day, the NO_{2} monitor at ground level was not operational. NO_{2} monitor data at 200 m were averaged over the lidar integration time. Bottom panel: Polluted day, September 12, 2006 (lidar measurement from 9:37–10:26 UT). For the NO_{2} monitor data at ground level an average was made for the time the monitor was operational during this interval, from 10:03–10:17. For the data of the NO_{2} monitor at 200 m two averages were determined; the lower average, for 9:37–10:17 UT, is for the situation that the NO_{2} monitor is above the boundary layer, the higher value, for 10:16–10:26 UT, is for the situation that the NO_{2} monitor is situated below the boundary layer.

D R A F T                   August 27, 2007, 12:40pm                   D R A F T

Figure 20. Two intex-b NO_{2} profiles. The upper panel shows data for a profile measured over the Mississippi-Alabama border (32.0 N, 88.3 W), while the lower panel shows data for a profile measured over the Gulf of Mexico (23.0 N, 91.1 W). Green dots are original measurements, the blue line is binned profile, and the red line is the annual mean geos-chem model profile for that location.

Figure 21. Tropospheric NO2 columns with in situ columns from INTEX-B. The symbols indicate land measurements (triangle) or ocean measurements (dot). The dotted line is a 1:1 ratio and the solid line is the fit to the data.

X - 66 CELARIER ET AL.: VALIDATION OF OMI NO_{2}

Table 3. Summary of validation study results for omi NO_{2} data product.

Instrument  Column Group  OMI, rel. to GB Remarks
saoz  Strat. CNRS  +10% As large as 50% in tropics.
max-doas  Trop. BIRA, etc.  -15% Large scatter in the data.
Brewer  Total GSFC  -33% Large N, large scatter.
mf-doas  Total WSU  -19% Very small N.
Pandora-1  Total GSFC  -15% Very small N.
ds-doas  Total BIRA  -16% Small N, large scatter.
ftuvs  Total JPL  -23% Small N, but good correlation.
Aircraft in situ Trop. UC Berkeley, GSFC  +10% Large scatter.